OpenStudy (anonymous):

Find the m<1. (Check my answer?) A.43 B.69 C.137 D.180

OpenStudy (anonymous):

I picked A. 

OpenStudy (anonymous):

103+34=137 180-137=43. Thats how i got my answer

OpenStudy (anonymous):

Is it right?

OpenStudy (anonymous):

well we know the angles of a triangle equal one 180 so what you did was find the angle inside the triangle next to 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

my answer is 69

OpenStudy (anonymous):

well we know the sum of the two interior angles equals the exterior of the opposite angle

OpenStudy (anonymous):

so really you had it right before with 103+34

OpenStudy (anonymous):

i think i know what i did wrong. my answer is 69 right?

OpenStudy (anonymous):

so my original answer is correct?

OpenStudy (anonymous):

which is it? youre confusing me..

OpenStudy (anonymous):

well one thing you know is that it is an exterior angle of a triangle so it has to be greater than 90

OpenStudy (anonymous):

answer my question i really dont have time to waste here.

OpenStudy (anonymous):

you just have to add the two interior angles of 103+34 to get the exterior angle of 1

OpenStudy (anonymous):

exterior angle 1*

OpenStudy (anonymous):

so 103+34=?